# Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech

### v1

@@ Abstract
Integrating human feedback to align text-to-speech (TTS) system outputs with human preferences has proven to be an effective approach for enhancing the robustness of language model-based TTS systems. Current approaches primarily focus on using preference data annotated at the utterance level. However, frequent issues that affect the listening experience often only arise in specific segments of audio samples, while other segments are well-generated. In this study, we propose a fine-grained preference optimization approach (FPO) to enhance the robustness of TTS systems. FPO focuses on addressing localized issues in generated samples rather than uniformly optimizing the entire utterance. Specifically, we first analyze the types of issues in generated samples, categorize them into two groups, and propose a selective training loss strategy to optimize preferences based on fine-grained labels for each issue type. Experimental results show that FPO enhances the robustness of zero-shot TTS systems by effectively addressing local issues, significantly reducing the bad case ratio, and improving intelligibility. Furthermore, FPO exhibits superior data efficiency compared with baseline systems, achieving similar performance with fewer training samples.
Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech
Jixun Yao1, Yuguang Yang2, Yu Pan2, Yuan Feng2, Ziqian Ning1, Jianhao Ye2, Hongbin Zhou2, Lei Xie1 1Northwestern Polytechnical University 2Everest Team, Ximalaya
1 Introduction
Large language models (LLMs) demonstrate remarkable in-context learning abilities in natural language processing (NLP) tasks by scaling both model parameters and dataset size (Brown et al., 2020; Touvron et al., 2023; Grattafiori et al., 2024). Reinforcement learning from human feedback (RLHF) (Christiano et al., 2017) is crucial for calibrating these models, aligning them with human preferences, and minimizing the risk of generating harmful or toxic content. This technique is essential for the success of systems like ChatGPT (Achiam et al., 2023) and has recently been extended to tasks in other modalities.
Speech, a key modality of generative artificial intelligence, is experiencing rapid advancements inspired by developments in text-based LLMs. Recent text-to-speech (TTS) studies (Wang et al., 2023; Chen et al., 2024b; Du et al., 2024) tokenize speech signals into discrete units and employ language model to generate speech tokens in a next-token prediction manner based on phoneme inputs and prompt speech recordings. By leveraging in-context learning capabilities, these TTS systems can produce high-quality, personalized speech in a zero-shot manner. Similar to the use of RLHF in calibrating text LLMs, there is growing interest in integrating RLHF into TTS systems (Zhang et al., 2024; Chen et al., 2024a; Hu et al., 2024; Tian et al., 2024). These approaches effectively enhance the robustness and zero-shot capabilities of TTS systems by incorporating human evaluation.
@@ 1 Introduction
Despite significant advancements in integrating RLHF into TTS systems, we believe two primary challenges remain. The first challenge is efficiently calibrating pre-trained TTS systems with limited preference data. RLHF typically requires preference data evaluated by human annotators, which is then used for training to align TTS systems with human preferences. However, obtaining large-scale human-annotated speech preference data is prohibitively expensive. Compared to annotating text data, annotating speech data is more difficult due to its inclusion of both linguistic content and paralinguistic information.
@@
The second challenge is achieving fine-grained alignment of specific segments in generated samples with human feedback. In recent years, issues such as poor speaker similarity and degraded audio quality, referred to as systemic errors, have been largely mitigated in TTS systems through advanced technology. However, more frequent issues that impact the listening experience often only occur in specific segments of the audio samples, which we call segmental errors. These include issues like mispronunciation of certain words, abnormal pauses, or repetition, etc. However, current speech RLHF approaches typically annotate preference data at the utterance level, which fails to account for fine-grained issues. Additionally, utterance-level preference optimization calculates the loss for the entire sequence, even though many tokens are already well-trained, leading to inefficiencies (Lin et al., 2024b). We believe that optimizing problematic segments in a fine-grained manner would be both more effective and more data-efficient.
@@
To address the above challenges, we propose a novel fine-grained preference optimization (FPO) framework—an RLHF-based approach aimed at addressing segmental errors in generated samples and improving the robustness of zero-shot TTS systems. FPO employs a fine-grained sampling-annotating pipeline, where representative audio samples are first generated multiple times by TTS systems. Preference data are then annotated from human listening perspectives, labeling each sample as preferred or dispreferred at the utterance level. For each paired preference data, we annotate the problematic segment of the local issue to obtain fine-grained token-level preference labels, enabling more precise optimization than traditional approaches. Furthermore, we analyze the types of issues across different problem segments from an evaluation perspective and propose a selective training loss strategy to optimize each type of issue. During training, FPO processes the entire sequence through the TTS models but computes loss exclusively for the problematic segments, which can effectively resolve fine-grained issues while enhancing training efficiency. Experimental results demonstrate that FPO significantly enhances the robustness of zero-shot TTS systems and exhibits superior data efficiency compared with baseline systems, improving the quality of generated speech across both subjective and objective metrics. The main contributions can be summarized as follows:
@@
  • We propose FPO, an RLHF-based approach designed to align problematic segments in generated samples, thereby enhancing the robustness of TTS systems. The key novelty of FPO lies in its preference optimization strategy, which focuses on optimizing specific problematic segments while avoiding unnecessary computation on well-learned tokens, resulting in a more data-efficient optimization process.
@@
  • By analyzing issues in generated samples, FPO fine-grained optimizes TTS systems on problematic segments using a selective training strategy. This approach introduces a new perspective for addressing fine-grained fixes in TTS preference optimization, focusing on resolving bad cases efficiently and effectively.
@@
  • Extensive experiments demonstrate the efficiency and effectiveness of FPO across both subjective and objective metrics. Notably, FPO significantly reduces the occurrence of bad cases compared to powerful baseline systems, effectively addressing the robustness challenges of zero-shot TTS systems. Audio samples can be found at https://yaoxunji.github.io/fpo.
2 Related Work
2.1 Neural codec based Zero-shot TTS
@@
Inspired by the success of LLMs in NLP, formatting the TTS task as a next-token prediction problem has gained significant popularity in recent years (Borsos et al., 2022; Zhang et al., 2023). VALL-E (Wang et al., 2023), as a pioneering work, introduced the first neural codec based TTS framework, leveraging large-scale speech data to develop in-context learning capabilities. It can synthesize high-quality personalized speech in a zero-shot manner using only a 3-second enrolled recording as an acoustic prompt. Building on this foundation, the typical neural codec-based TTS systems (Kharitonov et al., 2023; Xin et al., 2024; Peng et al., 2024) involve tokenizing speech signals into discrete tokens using a neural codec (Défossez et al., 2022; Zeghidour et al., 2021). A decoder-only language model is then trained to predict these discrete tokens alongside phoneme sequences. Finally, the predicted tokens are reconstructed into speech signals using the codec decoder. This approach has demonstrated promising scalability with large datasets and model sizes, enabling high-quality synthesized speech and powerful zero-shot capabilities for unseen speakers (Łajszczak et al., 2024; Du et al., 2024).
@@
2.2 Reinforcement Learning from Human Feedback
Human feedback has been widely used in language models for NLP tasks (Ouyang et al., 2022), with RLHF technology playing a crucial role in aligning models with human preferences. Conventionally, preference optimization is performed by maximizing rewards computed from a reward model to align generated content. Recent approaches, such as Direct Preference Optimization (DPO) (Rafailov et al., 2024), leverage closed-form loss functions that directly operate on preference data. This method eliminates the need for explicit reward model training while achieving alignment performance comparable to traditional RLHF.
@@ 2.2 Reinforcement Learning from Human Feedback
RLHF in TTS remains in its infancy compared to its advancements in NLP. SpeechAlign first introduced RLHF in TTS (Zhang et al., 2024), exploring multiple preference alignment methods using paired preference data. UNO (Chen et al., 2024a) and RIO (Hu et al., 2024) extended this by optimizing unpaired preference data: UNO accounts for annotation uncertainty in subjective evaluations, while RIO introduces a reverse preference data selection method based on Bayesian principles. Recent studies (Tian et al., 2024) have conducted thorough empirical evaluations of how preference alignment algorithms enhance LM-based TTS, and Seed-TTS (Anastassiou et al., 2024) also have adopted RLHF during the post-training stage of LM-based TTS. However, current approaches are limited to utterance-level preference optimization and overlook the potential of fine-grained alignment.
3 Methodology
3.1 Background
@@
Neural codec based zero-shot TTS comprises two components: a neural codec model and a language model. The codec model tokenizes the speech signal into discrete acoustic tokens $a\in\mathbb{R}^{n\times l}$, where $n$ is the number of quantizer layers in the codec model and $l$ is the temporal length. The tokenized acoustic tokens can later be reconstructed into speech signals by the codec decoder. Given a text transcript $t$, the language model predicts the acoustic tokens using a next-token prediction approach. A widely used codec model employs a single quantizer, resulting in the extracted acoustic token being a single sequence along the temporal dimension. This setup simplifies the process of directly concat $t$ and $a$ and computing their joint probabilities in an auto-regressive manner:
@@
RLHF with Preference Data. The language model is prompted with prompts $x$ to generate pairs of speech samples ($y_{1}$,$y_{2}$). Humans then evaluate these pairs and express preferences for one sample, denoted as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ represent the preferred and dispreferred samples respectively. The probability of preference data can be modeled using a reward model, which is trained via maximum likelihood estimation. During the reinforcement learning phase, the reward model provides feedback to the language model, which is then optimized by maximizing the rewards.
3.2 Preference Data Collection
The selection of preference data plays a critical role in RLHF. We propose a fine-grained sampling-annotation pipeline that begins with annotating pairwise preference data. Specifically, we first sample speech prompts from an unseen speaker pool for each text transcript and use them as input to the TTS system. To promote more diverse zero-shot TTS generation, we adjust sampling hyperparameters during inference and perform multiple inference runs for each sample, constructing a generated sample dataset $\{y_{1}^{b},y_{2}^{b},\ldots,y_{k}^{b}\}$, where $b$ is the $b$-th inference process.
After the sampling process, human annotators evaluate the generated samples from each inference process, identifying preferred and dispreferred ones. To mitigate the high human resource demands for annotations, we draw inspiration from previous studies (Lin et al., 2024a) and use pre-trained speech perceptual systems to simulate human feedback. We propose a comprehensive scoring method that combines multiple metrics—including intelligibility, quality, similarity, and duration—to determine the final score of each sample. This scoring system enables the identification of preferred samples based on their final scores $s(y)$:
@@ 3.2 Preference Data Collection
3.3 Fine-grained Annotation

### v2

# FPO: Fine-grained Preference Optimization Improves Zero-shot Text-to-Speech

@@ Abstract
Integrating reinforcement learning to align generated speech with human preferences has proven effective in improving the robustness of language model-based text-to-speech (TTS) systems. Current approaches primarily rely on preference data annotated at the utterance level. However, frequent issues affecting the listening experience, such as unnatural prosody, mispronunciations, or semantic inconsistencies, often arise only in specific segments of audio, while other segments may be well-generated and require no correction. This mismatch between coarse-grained supervision and fine-grained quality variation limits the effectiveness of preference-based optimization. In this study, we propose a fine-grained preference optimization approach (FPO) to enhance the robustness of TTS systems. FPO shifts the optimization paradigm from global utterance-level tuning to targeted local refinement, focusing on addressing localized issues in generated samples rather than uniformly optimizing the entire utterance. We begin by analyzing the types of common generation issues and categorizing them into temporal modeling errors and semantic-phonetic alignment errors, which frequently degrade intelligibility and naturalness. To tackle these problems, we introduce a selective training loss strategy that leverages fine-grained labels for each issue type, allowing the model to focus on learning signals where they are most needed. Experimental results demonstrate that FPO substantially improves the robustness of zero-shot TTS systems by effectively correcting problematic regions in the output. This leads to a significant reduction in the bad case ratio, improved intelligibility, and overall perceptual quality. Moreover, FPO exhibits strong data efficiency, achieving comparable or superior performance to baseline methods while requiring notably fewer training samples. Audio samples can be found at https://yaoxunji.github.io/fpo/.
I Introduction
Large language models (LLMs) demonstrate remarkable in-context learning abilities in natural language processing (NLP) tasks through scaling model parameters and dataset sizes [brown2020language, touvron2023llama, grattafiori2024llama3]. This scaling has enabled them to generalize across a wide range of tasks with minimal supervision, often achieving near-human or even superhuman performance in many domains. However, their vast and diverse training data encompasses varied goals, priorities, and content, not all of which are desirable to imitate. Selecting desired responses and behaviors from the model’s broad knowledge and capabilities is crucial for developing safe, high-performing, and controllable AI systems [christiano2017deep]. A key solution involves aligning LLMs with human preferences through reinforcement learning (RL), effectively calibrating outputs to reduce harmful content risks. This approach underpins widely-used systems like ChatGPT [achiam2023gpt] and has recently expanded beyond pure language generation to multimodal tasks, including vision-language understanding [kelly2024visiongpt], code generation [zheng2024opencodeinterpreter], and speech processing [yao2025gense].
Speech, a key modality in generative artificial intelligence, is experiencing rapid advancements inspired by developments in text-based LLMs. Recent text-to-speech (TTS) studies [wang2023valle, chen2024vall2, du2024cosyvoice] tokenize speech signals into discrete units via neural codec models, aligning them with phoneme or text inputs. These tokens are then processed by a language model to predict subsequent tokens in a next-token prediction manner using phoneme inputs and speech prompts. By scaling up model parameters and training data size, the in-context learning capabilities also emerge in TTS language models, enabling zero-shot generation of high-quality, personalized speech. Mirroring RL’s role in refining text LLMs, emerging interest focuses on applying RL to TTS systems [zhang2024speechalign, chen2024uno, hu2024rio, tian2024tx_rl]. Typical approaches introduce direct preference optimization (DPO) [rafailov2024dpo] or Kahneman-Tversky Optimization (KTO) [ethayarajh2024kto] to align TTS systems, while the key challenge is to collect desirable and relevant preference datasets. By annotating generated speech samples using professional annotators at the utterance-level, these approaches effectively enhance the speech quality and zero-shot capabilities of TTS systems by post-training in an RL manner.
@@ I Introduction
Despite significant advancements in integrating RL into TTS systems, we believe two primary challenges remain. The first challenge is efficiently calibrating pre-trained TTS systems with limited preference data. RL in TTS studies typically requires preference data evaluated by human annotators, which is then used to align TTS systems with human preferences. However, obtaining large-scale human-annotated speech preference data is prohibitively expensive. Compared to annotating text data, annotating speech data is more difficult due to its inclusion of both linguistic content and paralinguistic information.
@@
The second challenge is achieving fine-grained alignment of specific segments in generated samples with human feedback. While issues like poor speaker similarity and degraded audio quality—referred to as systemic errors—have largely been mitigated in TTS systems by scaling up training data, more frequent issues affecting the listening experience often occur in specific segments of the audio. We refer to these as segmental errors, which include mispronunciations, abnormal pauses, repetitions, and other disruptions. Current speech RL approaches typically annotate preference data at the utterance level, failing to capture these fine-grained issues. Moreover, using utterance-level preference data to optimize TTS systems needs to calculate losses of the entire sequence, even when many segments are already well-trained, leading to ineffective gradient updates and inefficient optimization [lin2024rho]. We believe that a more fine-grained approach to optimizing problematic segments would be both more effective and data-efficient.
@@
To address the above challenges, we propose a novel fine-grained preference optimization (FPO) framework for TTS systems—an RL-based approach designed to address segmental errors in generated samples and enhance the robustness of zero-shot TTS systems. FPO employs a fine-grained sampling-annotation pipeline for preference data collection, which plays a crucial role in the RL-based approaches. Pre-trained TTS systems first generate representative audio samples multiple times as the candidate preference data, while the utterance-level preference and dispreference data are annotated from human listening perspectives using these candidate data. For each paired preference data, we further identify the problematic segment and annotate it with fine-grained token-level preference labels, enabling more precise optimization than traditional methods. Furthermore, we analyze the types of issues across different segments from an evaluation perspective and divide the two types of errors in the annotated dispreference data. A selective training loss strategy is proposed to optimize different types of issues during preference optimization. Specifically, FPO processes the entire sequence through the TTS models but computes loss exclusively for the problematic segments, which can effectively resolve fine-grained issues while enhancing training efficiency. Experimental results demonstrate that FPO significantly enhances the robustness of zero-shot TTS systems and exhibits superior data efficiency compared to baseline systems, improving the quality of generated speech across both subjective and objective metrics. The main contributions can be summarized as follows:
@@
  • We propose FPO, an RL-based approach designed to align problematic segments in generated samples, thereby enhancing the robustness of TTS systems. The key novelty of FPO lies in its fine-grained preference optimization, which focuses on optimizing specific problematic segments rather than entire utterances. This approach avoids unnecessary computation on well-learned tokens, resulting in a more data-efficient optimization process.
@@
  • By analyzing issues in generated samples, FPO optimizes TTS systems by aligning problematic segments through a selective training strategy. This approach offers a new perspective for addressing fine-grained fixes in TTS preference optimization, which resolves bad cases both efficiently and effectively.
@@
  • Both subjective and objective experiments demonstrate the efficiency and effectiveness of FPO. Notably, FPO significantly reduces the occurrence of bad cases compared to baseline systems and can achieve similar performance with fewer training samples.
II Related Work
II-A Neural codec based Zero-shot TTS
@@
Inspired by the success of LLMs in NLP, formatting the TTS task as a next-token prediction problem has gained significant popularity in recent years [borsos2022audiolm, zhang2023speechgpt]. This formulation allows TTS systems to benefit from autoregressive modeling and in-context learning capabilities that have proven effective in text generation tasks. VALL-E [wang2023valle], as a pioneering work in this direction, introduced the first neural codec-based TTS framework, leveraging large-scale speech data to develop strong generalization and prompt-based adaptation abilities. It can synthesize high-quality, personalized speech in a zero-shot manner using only a 3-second enrolled recording as an acoustic prompt, without requiring speaker-specific fine-tuning or additional adaptation steps. Building on this foundation, a series of neural codec-based TTS systems [kharitonov2023speak, xin2024ralle, peng2024voicecraft] have been developed. These models typically involve a three-stage pipeline: (1) discretizing continuous speech waveforms into semantic and/or acoustic tokens using a neural audio codec [defossez2022high, zeghidour2021soundstream, zhang2023speechtokenizer, ji2024wavtokenizer], (2) training a decoder-only language model to predict these tokens conditioned on phoneme sequences and optional prompts, and (3) reconstructing the speech waveform from predicted tokens using the codec’s decoder. This discrete tokenization enables autoregressive modeling and makes the TTS task more analogous to language modeling.
Another prevalent approach introduces self-supervised speech representation models [hsu2021hubert, chen2022wavlm, chung2021w2v] to extract semantic tokens that capture linguistic content in a speaker-agnostic manner. These semantic tokens serve as intermediate targets for language models, which eases the token prediction task compared to directly modeling raw acoustic tokens[defossez2022high]. To reconstruct the waveform, many systems further employ diffusion-based vocoders or conditional decoders that synthesize speech from the generated semantic tokens, optionally guided by acoustic features or speaker embeddings. These neural codec-based frameworks have demonstrated strong scalability with increasing data and model size, enabling high-fidelity speech synthesis and robust zero-shot generalization to unseen speakers, speaking styles, and acoustic conditions [lajszczak2024base, du2024cosyvoice]. Despite their impressive performance, these models still face challenges in terms of robustness, particularly under mismatched conditions, and often suffer from localized quality degradations that are not effectively addressed by global training objectives.
@@
II-B Learning from Human Feedback
Human feedback has been widely utilized in LLMs for NLP tasks [ouyang2022training, stiennon2020learning], where RL techniques play a crucial role in aligning model behavior with human preferences. This alignment paradigm has been instrumental in developing AI agents that are helpful, harmless, and honest, paving the way for controllable language models. Traditionally, RL-based preference optimization follows a two-step pipeline. First, a reward model is trained on human-annotated preference data, where annotators compare model outputs and indicate which one is preferred. Then, the language model is fine-tuned using reinforcement learning, typically with Proximal Policy Optimization (PPO) [schulman2017proximal], to maximize the expected reward as estimated by the reward model. While effective, this approach is computationally intensive and sensitive to reward model quality and policy instability during RL fine-tuning.
To address these limitations, recent approaches such as Direct Preference Optimization (DPO) [rafailov2024dpo] propose a more efficient alternative. DPO introduces a closed-form loss function that operates directly on preference pairs, eliminating the need for an explicit reward model or reinforcement learning loop. This formulation simplifies the training pipeline while achieving alignment performance comparable to traditional RLHF methods. Other variants [gu2025mask], such as Implicit Preference Optimization (IPO) [garg2025ipo] and Kahneman-Tversky Optimization (KTO) [ethayarajh2024kto], explore related directions to improve preference learning stability and generalization. Although these methods have been extensively applied to NLP tasks, their adoption in speech-related domains remains relatively limited. Recent works begin exploring learning from human preferences in speech synthesis and voice conversion, yet most rely on utterance-level binary labels, which can be insufficient to capture the fine-grained nature of perceptual quality issues in speech.
@@ II-B Learning from Human Feedback
Learning from human feedback in TTS remains in its infancy compared to its advancements in NLP. SpeechAlign first introduced RLHF in TTS [zhang2024speechalign], exploring multiple preference alignment methods using paired preference data. UNO [chen2024uno] and RIO [hu2024rio] extended this by optimizing unpaired preference data: UNO accounts for annotation uncertainty in subjective evaluations, while RIO introduces a reverse preference data selection method based on Bayesian principles. Recent studies [tian2024tx_rl] have conducted thorough empirical evaluations of how preference alignment algorithms enhance LM-based TTS, and Seed-TTS [anastassiou2024seedtts] has also adopted the RL method during the post-training stage of LM-based TTS. However, current RL-based speech synthesis approaches are limited to utterance-level preference optimization and overlook the potential of fine-grained alignment. While token-level alignment methods (e.g., Mask-DPO [gu2025maskdpo], Rho-1 [lin2024rho]) have been explored in NLP, we adapt and extend it to speech-specific issues such as temporal modeling and semantic-phonetic misalignment, which are absent in text-based tasks.
III Methodology
III-A Background
@@
Neural codec based zero-shot TTS comprises two components: a neural codec model and a language model. The codec model tokenizes the speech signal into discrete acoustic tokens $a\in\mathbb{R}^{n\times l}$, where $n$ is the number of quantizer layers in the codec model and $l$ is the temporal length. The tokenized acoustic tokens can later be reconstructed into speech signals by the codec decoder. Given a text transcript $t$, the language model predicts the acoustic tokens using a next-token prediction approach. A widely used codec model employs a single quantizer [ye2025llasa], resulting in the extracted acoustic token being a single sequence along the temporal dimension. This setup simplifies the process of directly concat $t$ and $a$ and computing their joint probabilities in an auto-regressive manner:
@@
RL with Preference Data. The language model is first employed to generate paired speech samples ($y_{1}$$y_{2}$) with prompts $x$. Humans then evaluate these pairs and express preferences for one sample, denoted as $y_{w}\succ y_{l}\mid x$, where $y_{w}$ and $y_{l}$ represent the preferred and dispreferred samples respectively. The probability of preference data can be modeled using a reward model, which is trained via maximum likelihood estimation. During the reinforcement learning phase, the reward model provides feedback to the language model, which is then optimized by maximizing the rewards.
III-B Preference Data Collection
The selection of preference data plays a critical role in the RL process. We propose a fine-grained sampling-annotation pipeline that begins with annotating pairwise preference data. Specifically, we first sample speech prompts from an unseen speaker pool for each text transcript and use them as input to the TTS system. To promote more diverse zero-shot TTS generation, we adjust sampling hyperparameters during inference and perform multiple inference times for each prompt, constructing a generated sample dataset $\{y_{1}^{b},y_{2}^{b},\ldots,y_{k}^{b}\}$, where $b$ is the $b$-th inference process.
After the sampling process, human annotators evaluate the generated samples from each inference process, identifying preferred and dispreferred ones. To mitigate the high human resource demands for annotations, we draw inspiration from previous studies [lin2024alignslm] and use pre-trained speech perceptual systems to simulate human feedback. We propose a comprehensive scoring method that combines multiple metrics—including intelligibility, quality, similarity, and duration—to determine the final score of each sample. This scoring system can identify preferred samples based on their final scores $s(y)$:
@@ III-B Preference Data Collection
III-C Fine-grained Annotation
To achieve fine-grained alignment of TTS systems, we first thoroughly investigate the characteristics of segmental errors and categorize them into two types: temporal modeling errors and semantic-phonetic alignment errors. As illustrated in Figure 1, temporal modeling errors stem from the model’s limitations in capturing the temporal dynamics of speech, leading to issues such as mispronunciation, where abrupt or unstable pronunciation changes occur due to insufficient modeling of pitch continuity or variation trends. Another common problem is abnormal silence, where the model erroneously inserts silent frames or extends silent segments, disrupting the naturalness of generated speech. Additionally, the generated speech may exhibit unnatural prosody (i.g. unnatural pause), as the model may fail to adequately capture local rhythmic patterns, including stress, duration, and speaking rate, resulting in speech that deviates from natural prosody.